### v2

# Consistencies and inconsistencies in redshift-independent distances

Abstract
Redshift-independent distances underpin much of astrophysics, and there exists a plethora of methods to estimate them. However, the extent to which the distances they imply are consistent, while crucial for the integrity of the distance ladder, has been little explored. We construct a statistical framework to assess both internal (between measurements with the same method) and external (between-method) consistency by comparing differences between distances to their quoted statistical uncertainties in the NASA/IPAC Extragalactic Database of Distances (NED-D). 66 of the 76 indicators in NED-D are amenable to a consistency test by having at least two measurements to the same galaxy or at least one measurement to a galaxy also measured by another method. We find that only 12 of these methods produce self-consistent distances across literature determinations, of which 7 are also consistent with distances to the same galaxies measured by all other methods. The most consistent 6 methods (M-stars luminosity, Novae, Masers, Globular Cluster Fundamental Plane, O- and B-type Supergiants and BL Lac Luminosity) also give similar average distances to the mean of all indicators, while the 7th (Proper Motion) underestimates distances relative to the mean by 17.1%. We also investigate consistency of Cepheid distances in the SH0ES 2022 catalogue, finding no evidence for unaccounted-for systematics. Our NED-D results imply that considerable work remains to obtain reliable distances by a multitude of methods, a crucial endeavour for constructing a multiply cross-checked and fully robust distance ladder. We make our code publicly available.
keywords:
galaxies: distances and redshifts – cosmology: distance scale – methods: statistical – astronomical data bases
pubyear: 2025pagerange: Consistencies and inconsistencies in redshift-independent distancesLABEL:LastPage
1 Introduction
[…]
Assuming a cosmological model (e.g. $\Lambda$CDM) with calibrated values for its free parameters, distances in the Universe may be estimated from objects’ redshifts alone. However, within the local $\sim$100 Mpc redshifts provide inaccurate distance estimates because galaxies’ velocities are not predominantly due to the expansion of the Universe. There is a significant contribution from “peculiar velocities”, galaxies’ individual motions superimposed on the Hubble flow generated by gravitational interaction with their cosmic environment (for reviews see Strauss & Willick 1995; Turner 2024). If distances to such galaxies can be measured redshift-independently then peculiar and expansion velocities can be determined separately. The pattern of peculiar velocities provides evidence concerning the large-scale structure of the local Universe (Dupuy & Courtois, 2023; Valade et al., 2024), the growth rate of cosmic structure (Boruah et al., 2020; Said et al., 2020; Stiskalek et al., 2025), extensions to General Relativity (Lyall et al., 2023, 2024), the large-scale isotropy of the Universe (e.g. Feindt et al. 2013; Watkins et al. 2023) and the typicality or otherwise of the Milky Way’s place in the Universe (Bovy et al., 2009; McMillan, 2016). It also allows the distance–redshift relation to test, rather than assume, cosmological models.
Given these many and varied opportunities, much effort has naturally been devoted to measuring redshift-independent (or direct) distances. This is done in a series of steps, likened to rungs of a ladder: direct geometric methods (mainly parallax) are used to establish distances to nearby objects (mainly stars within the Milky Way), which are in turn used to calibrate objects on the “first rung” of the distance ladder. The most common such objects are Cepheid stars—which obey a linear relation between their distance-independent pulsation period and distance-dependent luminosity (for reviews see Freedman & Madore 2024; Anderson 2024)—and the Tip-of-the-Red-Giant-Branch (TRGB) feature of the colour–magnitude diagram, which has a fixed luminosity arising from the Helium flash in an electron-degenerate stellar core of intermediate-to-low-mass stars. Hence, it acts as a standard candle (Lee et al. 1993; for a review see Li & Beaton 2024). Once calibrated, these methods can be used to establish direct distances to $\sim$50 Mpc. They in turn are used to calibrate the ladder’s “second rung”. A key method here is SNIa which are standardisable candles: after correction for the colour and width of their light-curves they have a fixed luminosity, so that if this is calibrated using e.g. Cepheids or TRGB the measured flux determines the distance. This allows the Hubble diagram to be extended out to $z\approx 1$ ($\sim$3 Gpc; Brout et al. 2022; Rubin et al. 2023; Vincenzi et al. 2024).
Parallax–Cepheid/TRGB–SNIa is the best-known route to the distance ladder, but there is in fact a plethora of methods for determining redshift-independent distances. Based on standard candle, standard ruler, standard siren and other techniques, these employ different objects, different physics, different measurement methods and different systematics, and operate at various locations of the ladder (i.e. requiring varying degrees of calibration). Given the crucial importance of the distance ladder for cosmology, and the tensions it has thrown up, it seems prudent to compare carefully as wide as possible a range of methods to assess possible systematic errors and with an eye to swapping out or supplementing some methods for others in constructing a maximally robust, precise and far-reaching ladder.
That is the aim of this paper. We take as our starting point the NASA/IPAC Extragalactic Database of Distances (NED-D), [^1] https://ned.ipac.caltech.edu/Library/Distances/ a compilation containing $>$270,000 extragalactic distances measured to $>$140,000 galaxies using 76 different distance indicators (Steer et al., 2016). Our particular focus here is on quantifying the statistical consistency of distances measured to the same galaxies, either of multiple estimates with the same method (self-consistency) or between methods (cross-consistency). This depends crucially on the quoted uncertainties, which are themselves model-dependent quantities with their own uncertainties. We apply the same methods to the Supernova $H_{0}$ for the Equation of State (SH0ES) pipeline as a leading example of distance ladder-based $H_{0}$ inference framework (Riess et al., 2009, 2011, 2012, 2022). Our approach will allow us to determine which methods appear most reliable, and whether systematics may be biasing literature inferences employing the distance ladder. In case of inconsistency we can establish whether this is due to under- or over-estimated uncertainties and/or mean shifts in all distances inferred by some method. This provides a stepping stone towards construction of a more confidently reliable ladder employing multiple redshift-independent distance methods.
Our work builds on that of Singh (2021), who identified a systematic underestimation in uncertainties in NED-D that is most prominent for the most widespread methods and growing over time. This suggests that the precision of e.g. $H_{0}$ inferred from the distance ladder may be overstated, clearly of great importance to the aforementioned tension. Our improvements on this analysis are first to improve the statistical measure and test of consistency and second to perform an independent analysis of the SH0ES 2022 results (Riess et al., 2022) as the first catalogue claiming a $5\sigma$ Hubble tension. This provides information on whether inconsistencies in distance measurements may be responsible for (some of) the tension.
The structure of the paper is as follows. In Sec. 2 we describe the NED-D and SH0ES data that we analyse. Sec. 3 details our statistical methodology, while Sec. 4 presents our findings. We discuss more general ramifications of the work as well as caveats in Sec. 5, and conclude in Sec. 6.
2 Observational Data
2.1 NED-D
NED-D is an online compilation of redshift-independent extragalactic distances containing $>270,000$ distance measurements to $>140,000$ galaxies using 76 different methods (Steer et al., 2016). These methods, categorised as standard candles, standard rulers or “secondary” (defined as having a precision $\sim$20as opposed to the other “primary” methods with precision $\lesssim$10%; Steer et al. 2016), are as follows:
  • Standard candles: AGN time lag, Asymptotic Giant Branch Stars (AGB), B-type Stars (B Stars), BL Lac Object Luminosity (BL Lac Luminosity), Black Hole, Blue Supergiant, Brightest Cluster Galaxy (BCG), Brightest Stars, Carbon Stars, Cepheids, Colour-Magnitude Diagrams (CMD), Delta Scuti, Flux-Weighted Gravity-Luminosity Relation (FGLR), Gamma-Ray Burst (GRB), Globular Cluster Luminosity Function (GCLF), Globular Cluster Surface Brightness Fluctuations (GC SBF), HII Luminosity Function (HII LF), Horizontal Branch, M Stars luminosity (M Stars), Miras, Novae, O- and B-type Supergiants (OB Stars), Planetary Nebula Luminosity Function (PNLF), Post-Asymptotic Giant Branch Stars (PAGB Stars), Quasar spectrum, RR Lyrae Stars, Red Clump, Red Supergiant Variables (RSV Stars), Red Variable Stars (RV Stars), S Doradus Stars, SNIa SDSS, SX Phoenicis Stars, Short Gamma-Ray Bursts (SGRB), Statistical, Subdwarf Fitting, Sunyaev-Zeldovich Effect (SZ effect), Surface Brightness Fluctuations (SBF), Tip of the Red Giant Branch (TRGB), Type II Cepheids, Type II Supernovae Radio (SNII radio), Type Ia Supernovae (SNIa), White Dwarfs, Wolf-Rayet, Gravitational Wave (Grav. Wave.)
  • Standard rulers: CO ring diameter, Dwarf Galaxy Diameter, Eclipsing Binary, Globular Cluster Radii (GC radius), Grav. Stability Gas. Disk, Gravitational Lenses (G Lens), HII Region Diameters (HII), Jet Proper Motion, Masers, Orbital Mechanics (Orbital Mech.), Proper Motion, Ring Diameter, Type II Supernovae, Optical (SNII Optical).
  • Secondary methods: D-Sigma, Diameter, Dwarf Ellipticals, Faber-Jackson, Fundamental Plane (FP), GC K vs. (J-K), GeV TeV ratio, Globular Cluster Fundamental Plane (GC FP), H I + optical distribution, Infra-Red Astronomical Satellite (IRAS), L(Hbeta)-sigma, LSB galaxies, Magnitude, Mass Model, Radio Brightness, Sosies, Tertiary, Tully Estimate (Tully est), Tully-Fisher.
[…]
2.2 SH0ES 2022
We also consider the Cepheid catalogue used by the SH0ES collaboration in 2022. This is the latest full release of the SH0ES data. Furthermore, this is particularly pertinent as it has been used to argue for a $>5\sigma$ Hubble tension (Riess et al., 2022).
The SH0ES 2022 catalogue includes 3130 Cepheid measurements across 40 galaxies. Two of them correspond to first-rung galaxies with direct geometric distances, one non-SNIa host galaxy, with the remainder in the second rung used to calibrate SNIa. The third calibration galaxy, the Milky Way (MW), imposes an effective Gaussian prior on the Wesenheit absolute magnitude at a period of 10 days and solar metallicity $M^{W}_{H,1}$. The full catalogue is available in three FITS files. The data does not include the distance modulus $\mu$ to each Cepheid, but rather the period $P$, Wesenheit apparent magnitude $m^{W}_{H}$ in the F160W passband and excess metallicity $[O/H]$ along with their covariance matrix. The absolute magnitude can be derived from the Period–Luminosity (PL) relation via (Riess et al., 2022)
$\mu_{0}=m^{W}_{H}-M^{W}_{H,1}-b_{W}(\log_{10}P-1)-Z_{W}[O/H],$ (1)
where $M^{W}_{H,1}$ is the fiducial Wesenheit absolute magnitude at $P=10$ days and solar metallicity, and $b_{W}$ and $Z_{W}$ are parameters defining the empirical relation between these variables.
Riess et al. (2022) quote constraints for these parameters, which would enable us to compute the individual distance modulus for all Cepheids in the catalogue. However, it does not include the full covariance matrix necessary to derive their uncertainties. Thus, we ran a Monte Carlo Markov Chain (MCMC) analysis using the emcee sampler [^4] https://emcee.readthedocs.io/en/stable/ (Foreman-Mackey et al., 2013) with 100 walkers and $>50$ times the autocorrelation length in samples to ensure convergence. The set of inferred parameters includes the distance modulus to 38 galaxies, the differences between the predicted and geometric distance moduli for the two first-rung galaxies (minimised to calibrate the PL relation), the PL relation parameters previously mentioned, the absolute magnitude of SNIa $M^{0}_{B}$, $\Delta zp$ (a parameter accounting for possible systematics associated with ground-based observations), and $5\log H_{0}$.
[…]
$\text{Cov}(\mu_{0k},\mu_{0l})=\sum_{i=1}^{M}\sum_{j=1}^{M}\left(\dfrac{\partial\mu_{0k}}{\partial\theta_{i}}\right)\left(\dfrac{\partial\mu_{0l}}{\partial\theta_{j}}\right)\text{Cov}(\theta_{i},\theta_{j}),$ (2)
where $\theta_{i,j}$ are the parameters $\{m_{H,(k,l)}^{W},M_{H,1}^{W},b_{W},Z_{W}\}$ ($M=4$). This ensures that the uncertainties are properly propagated into the distance moduli. With this result we can perform our consistency analysis on this catalogue. We carry this out on the whole catalogue and also on each galaxy separately to search for galaxy-specific systematics.
3 Methodology
3.1 Quantifying consistency
We perform two kinds of consistency tests: internal (or self) and external (or cross). For the internal test we consider the measurements for a given method and a given galaxy with each other, while for the external test we compare for each galaxy the measurements from every pair of methods. This naturally restricts us to galaxies for which there are multiple measurements.
We compute the difference between each pair of measurements using
$\Delta_{ij}=\sqrt{(\mu_{i}-\mu_{j})^{T}\>\Sigma_{ij}^{-1}\>(\mu_{i}-\mu_{j})}\>,$ (3)
where $\mu_{i,j}$ are the distance moduli, and $\Sigma$ is the covariance between the two measurements. In this case $\Sigma_{ij}$ is a number given by $\sigma_{\mu_{i}}^{2}+\sigma_{\mu_{j}}^{2}-2\,\text{Cov}(\mu_{i},\mu_{j})$ where $\sigma_{\mu_{i,j}}$ are the diagonal uncertainty terms and $\text{Cov}(\mu_{i},\mu_{j})$ the cross-term between the two measurements. Note that $\text{Cov}(\mu_{i},\mu_{j})$ is assumed to be zero for all NED-D measurements, and hence only plays a role in the SH0ES analysis. Eq. 3 gives the Mahalanobis distance (Mai-Jalanobis, 1936) describing the difference in units of $\sigma$ between the measurements. It would be distributed as a standard half-normal if $\mu_{i}$ and $\mu_{j}$ were scattered from a common true value by their uncertainties (positive part of the standard normal $\mathcal{N}(\mu=0,\sigma^{2}=1)$ for $x\geq 0$). [^5] Eq. 3 may be compared to eq. 1 of Singh (2021), $\Delta_{12}=|\mu_{1}-\mu_{2}|/\sigma_{\mu_{1}}$). This is not symmetric under $\mu_{1}\leftrightarrow\mu_{2}$, although the combination effectively used in that work, $(\Delta_{12}+\Delta_{21})/2$, is. Our formula has the justification of being derived from the convolution of two Gaussians assumed to be independent.
We calculate the following quantities:
  1. 1.
    For each method in the NED-D catalogue, we take all the galaxies having at least two measurements. We compute $\Delta_{ij}$ for all possible pairs across all galaxies in that method. This quantifies internal consistency.
  2. 2.
    For all possible pairs of methods in the NED-D catalogue, we take all galaxies having at least one measurement by both methods. We compute $\Delta_{ij}$ for all possible pairs across all galaxies for this method pair. This quantifies external consistency.
  3. 3.
    For the SH0ES 2022 catalogue, we compute $\Delta_{ij}$ for all possible measurement pairs across all galaxies. This is a global measure of consistency across the catalogue.
  4. 4.
    For each galaxy individually in the SH0ES 2022 catalogue, we compute $\Delta_{ij}$ for all possible measurement pairs in this galaxy. This measures consistency for each galaxy separately.
This produces in each case a $\Delta_{ij}$ distribution, which we then compare to the expected standard half-normal distribution. If the $\Delta_{ij}$ could not plausibly have been drawn from this distribution this implies an inconsistency, which could be due to unaccounted-for systematic errors, underestimated or overestimated uncertainties, outliers, inconsistent normalisations of distance and/or other problems. To assess this consistency we compare the $\Delta_{ij}$ and standard half-normal distributions using the Kullback–Leibler divergence (KLD; Kullback & Leibler 1951; MacKay 2003). [^6] One might think first to use the Kolmogorov–Smirnov (KS) test (An, 1933; Smirnov, 1948). This is however unsuitable, as the $p$-value of a given test statistic depends on the size of the input array, here the number of measurement pairs. That this is not the number of independent measurements biases $p$ low, which can lead one to infer inconsistency even for mock data drawn from a standard half-normal when the sample size is large. This can be addressed in the case of self-consistency by computing the $p$-value using the number of independent measurements rather than the number of pairs, but it is not possible to assign unambiguously a number of effective samples in the cross-consistency test. This is defined by
[…]
where $P$ is the empirical probability distribution function (pdf), $Q$ is the comparison pdf (standard half-normal). We validated our results by computing the KL divergence using the Simpson’s rule and the entropy (given by $\sum_{i}P(x_{i})\log(P(x_{i})/Q(x_{i}))$) and found that they give consistent comparable results with each other. Eq. 4 measures the similarity of the two pdfs. To convert the discrete measurements into a continuous KL pdf we apply Gaussian Kernel Density Estimation (KDE; Virtanen et al. 2020) with a bandwidth of the form $h=Cn^{-1/5}$ where $n$ is the number of measurement pairs. We then apply either Scott’s (Scott, 2015) or Silverman’s (Silverman, 2018) method for setting the bandwidth, which we find to give near-identical results. We also find similar results using a histogram (top-hat KDE) rather than Gaussian KDE.
This measures the “distance” between the measured distribution and standard half-normal, but does not by itself tell us how unexpected any such distance is. To determine this, we compute the KLD of 10,000 mock datasets generated assuming that the $\Delta_{ij}$ test statistic precisely follows a standard half-normal. This produces the KLD distribution corresponding to the null hypothesis that the measurements are perfectly consistent up to scatter described by their uncertainties. This allows us to compute a $p$-value for rejection of the null hypothesis as the fraction of mock datasets having a KLD greater than the empirical one. We use a threshold of 0.05 to reject the null hypothesis. If $p>0.05$ there is insufficient evidence to conclude that the measurements are not simply scattered by their uncertainty distributions; we term this “consistency”. For the simulations to be on equal footing to the real data, we simulate the same number of measurement pairs as exists in NED-D. This ensures that the $p$-value is independent of the number of galaxies and measurements.
Our choice to define the null hypothesis as consistency between the measurements means that significant evidence is required before concluding inconsistency. One could imagine instead defining the null hypothesis as inconsistency, so that significant evidence would be required to conclude consistency. Our choice has two advantages. First, while “consistency” is a well-defined model (each galaxy has an underlying true distance, which produces the various measurements through scattering by the observational uncertainties), inconsistency is not: there is an infinity of ways in which two sets of measurements may be inconsistent. It is therefore difficult to see how a null hypothesis of inconsistency could be unambiguously defined. Second, our choice makes our results conservative in the sense that the inconsistencies that we identify are significant by construction. We caution however that while we apply the term “consistency” to any comparison yielding $p>0.05$, we really mean only that the null hypothesis of consistency cannot be robustly ruled out. A $p$-value greater than 0.05 may still favour inconsistency. In some sense our results therefore provide a lower bound on the degree of inconsistency in the datasets we analyse.
Note that defining the $p$-value as the fraction greater than the measured value corresponds to us doing a one-tailed hypothesis test, searching for inconsistency only in the direction that the real data is less like a standard normal than the mocks. It is in fact possible for the data to be abnormal in the other direction, viz. more like a standard normal than would be expected under the null hypothesis. This would be tested by a two-tailed test that calculates $p$ as the fraction of mock datasets more extreme than the real data. We choose simply to flag such “anomalously good” cases, which correspond to the one-tailed $p$-value being $>0.975$.
3.2 Extracting insight from inconsistency
The result of Sec. 3.1 is a $p$-value for rejection of the null hypothesis of consistency between each method and any other (including itself) in the case of NED-D, and between all the Cepheid distances in the SH0ES catalogue either globally or galaxy-by-galaxy.
There are a number of insights we can draw from this. The first is to identify the “best” (most consistent) distance indicators in the NED-D catalogue. To achieve this, we order the methods in increasing external consistency fraction. For a given indicator, we consider all other indicators that share at least one galaxy. We define the external consistency fraction as the number of such pairs that are consistent divided by the total number. Then, we remove the method with the lowest percentage external consistency and again compute the external consistency percentage assuming that a discarded method is responsible for inconsistency in any pair containing it. We repeat this process until we remain with a set of methods that have complete external consistency with each other. This is the maximal set of externally consistent methods.
Next we can learn about the cause of the inconsistencies. To learn whether an inconsistency derives under- or overestimated uncertainties we perform an additional analysis by rescaling the distance uncertainties by the ratio of the uncertainty of each galaxy to the standard half-normal standard deviation ($\sigma^{\prime}_{i}=(\Sigma/\sqrt{1-2/\pi})\sigma_{i}$ where $\sigma^{\prime}_{i}$ is the rescaled uncertainty of a given distance measurement, $\Sigma$ is the standard deviation of the empirical pdf before the rescaling and $\sigma_{i}$ is the original uncertainty). If this produces consistency, the original inconsistency must have been caused by the uncertainties. In that case, we then compare the variance of the $\Delta_{ij}$ distribution to that of the standard half-normal (1-2/$\pi$). If it is higher, then the uncertainties are underestimated, and vice versa.
For the case of external consistency the situation is more complex as an inconsistency could derive from inconsistent means of the $\mu_{i}$ and $\mu_{j}$ distributions rather than anything to do with the uncertainties. To disentangle the possible causes, we repeat the consistency analysis with two modifications. First, for each galaxy in the method pair, we computed the mean distance modulus over the measurements of each method. We then shift the mean of one method onto that of the other and repeat the consistency test. If this transformation produces consistency, the original inconsistency must have been due to inconsistent means. The second modification is instead of shifting the means to rescale the uncertainties by the relative standard deviations of the $\mu$ distributions. If this produces consistency, the original inconsistency must have come from the uncertainties. Note that having inconsistent means and uncertainties are not mutually exclusive. In some cases consistency may be achieved either by shifting the means or rescaling the uncertainties; in others doing both is required. If the test returns inconsistency even after both of these are done, it must be due to other effects such as outliers, non-Gaussian uncertainties or systematic errors. We call this an “essential inconsistency,” indicating that it cannot (easily) be removed.
Table 1: Results for the standard candle distance indicators. We present the number of galaxies having at least two measurements by a given method, the number of measurement pairs in these galaxies, the total number of measurements in NED-D along with the number of papers from which these measurements derive, the KLD $p$-value and external consistency statistics. The $p$-value is shown in green if it is greater than 0.05 (indicating consistency, i.e. not enough evidence to reject the null hypothesis), orange if it is between 0 and 0.05 (indicating inconsistency), and red if it is 0 (indicating severe inconsistency). In cases of inconsistency, we indicate the $p$-value after rescaling the uncertainties to a standard half-normal in parenthesis. If this produces consistency, we indicate whether the uncertainties were originally underestimated (downward-pointing arrow) or overestimated (upward-pointing arrow). The final column shows the fraction of other methods with which the method in question is externally consistent: the first number gives the number of cases that are consistent out-of-the-box, the second the number that are made consistent by shifting the means but not by rescaling the uncertainties, the third the number that are made consistent by rescaling the uncertainties but not by shifting the means, the fourth the number that are made consistent by either shifting the means or rescaling the uncertainties, and the fifth the number that are made consistent only by both shifting the means and rescaling the uncertainties. The final number shows the total number of other methods with which a given method is compared, such that this number minus the sum of the others is the number of essential inconsistencies.
Methods that are internally consistent and members of the maximal externally consistent set are shown in bold. (See text for full details.)
Distance indicator # of galaxies # pairs # total measurements/references KL $p$-value External consistency
BL Lac Luminosity 2 4 20 / 11 0.9705 1 - 0 - 0 - 0 - 0 / 1
RSV Stars 2 4 9 / 6 0.8298 9 - 0 - 3 - 4 - 0 / 28
OB Stars 2 2 4 / 3 0.7383 27 - 0 - 1 - 0 - 0 / 31
Novae 6 14 18 / 11 0.4382 35 - 0 - 0 - 3 - 1 / 46
Miras 6 119 36 / 27 0.1121 20 - 0 - 4 - 2 - 0 / 39
RV Stars 1 10 5 / 1 0.0845 6 - 0 - 1 - 0 - 0 / 18
Statistical 14 168 292 / 27 0.0795 15 - 0 - 10 - 3 - 0 / 54
Brightest Stars 21 127 129 / 41 0.074 21 - 0 - 6 - 3 - 2 / 44
M Stars 1 3 6 / 5 0.0686 18 - 0 - 1 - 1 - 0 / 33
Delta Scuti 2 29 12 / 2 0.0395 (0.8241) $\uparrow$ 14 - 0 - 4 - 1 - 0 / 31
Horizontal Branch 22 139 110 / 53 0.0037 (0.1495) $\downarrow$ 19 - 0 - 4 - 4 - 0 / 40
PNLF 45 285 183 / 57 0.0016 (0.0019) 23 - 0 - 12 - 3 - 1 / 52
AGN time lag 11 13 32 / 3 0.0 (0.7099) $\downarrow$ 2 - 0 - 3 - 1 - 0 / 6
Carbon Stars 7 182 82 / 13 0.0 (0.4555) $\uparrow$ 25 - 0 - 1 - 1 - 0 / 38
Cepheids 115 69496 10956 / 325 0.0 (0.0) 8 - 0 - 6 - 2 - 1 / 47
CMD 123 41733 1265 / 170 0.0 (0.0) 15 - 0 - 6 - 5 - 1 / 51
GRB 139 1661 729 / 17 0.0 (0.0) -
GCLF 176 2288 781 / 96 0.0 (0.0) 15 - 0 - 4 - 1 - 0 / 40
RR Lyrae 2443 16262 26409 / 216 0.0 (0.0) 16 - 1 - 4 - 2 - 2 / 43
Red Clump 24 3420 236 / 63 0.0 (0.0) 16 - 0 - 4 - 1 - 1 / 40
SNIa SDSS 1849 7687 6054 / 1 0.0 (0.0) 0 - 0 - 2 - 0 - 0 / 4
SZ effect 78 426 283 / 19 0.0 (0.0) 6 - 0 - 0 - 1 - 0 / 16
SBF 406 4403 1781 / 67 0.0 (0.0) 19 - 0 - 8 - 2 - 1 / 42
TRGB 307 8537 1736 / 336 0.0 (0.0) 15 - 1 - 3 - 9 - 0 / 54
Type II Cepheids 7 435 62 / 22 0.0 (0.3613) $\downarrow$ 18 - 1 - 6 - 1 - 0 / 44
SNIa 4978 129581 30843 / 140 0.0 (0.0) 10 - 0 - 5 - 2 - 0 / 38
AGB 0 0 2 / 2 - 22 - 0 - 0 - 0 - 0 / 23
B Stars 0 0 2 / 2 - 28 - 0 - 0 - 0 - 0 / 30
Black Hole 0 0 17 / 2 - 4 - 1 - 2 - 1 - 0 / 11
Blue Supergiant 0 0 2 / 1 - 11 - 0 - 0 - 0 - 0 / 25
BCG 0 0 239 / 1 - 9 - 0 - 2 - 0 - 0 / 11
FGLR 0 0 7 / 7 - 9 - 0 - 3 - 4 - 0 / 21
GC SBF 0 0 3 / 1 - 13 - 0 - 2 - 0 - 0 / 19
HII LF 0 0 17 / 2 - 6 - 0 - 0 - 1 - 0 / 15
SX Phe Stars 0 0 3 / 2 - 10 - 0 - 0 - 0 - 0 / 13
Subdwarf fitting 0 0 1 / 1 - 19 - 0 - 0 - 0 - 0 / 22
SNII radio 0 0 24 / 2 - 3 - 1 - 6 - 4 - 0 / 17
White Dwarfs 0 0 1 / 1 - 21 - 0 - 0 - 0 - 0 / 22
Wolf-Rayet 0 0 1 / 1 - 0 - 0 - 1 - 3 - 0 / 5
Grav. Wave 0 0 8 / 8 - 4 - 0 - 0 - 0 - 0 / 4
Table 2: As Table 1, but for standard ruler (above the horizontal line) and secondary (below the line) distance methods.
Distance indicator # of galaxies # pairs # total measurements/references KL $p$-value External consistency
Proper Motion 1 6 5 / 3 0.8918 24 - 2 - 3 - 0 - 0 / 32
Maser 6 29 24 / 20 0.6253 28 - 0 - 4 - 2 - 0 / 36
Dwarf Galaxy Diameter 8 8 16 / 1 0.0004 (0.582) $\uparrow$ 17 - 0 - 3 - 0 - 0 / 21
G Lens 15 70 86 / 14 0.0002 (0.6975) $\downarrow$ 1 - 0 - 2 - 0 - 0 / 3
Eclipsing Binary 49 3397 273 / 48 0.0 (0.0007) 22 - 0 - 2 - 1 - 0 / 42
HII region diameter 19 25 53 / 7 0.0 (0.0261) 7 - 0 - 15 - 7 - 0 / 42
Ring Diameter 47 47 212 / 1 0.0 (0.2025) $\uparrow$ 10 - 0 - 6 - 1 - 0 / 18
SNII optical 309 6090 1622 / 89 0.0 (0.0) 17 - 0 - 4 - 3 - 0 / 42
GC radius 0 0 101 / 8 - 22 - 0 - 9 - 2 - 0 / 37
Grav. Stability Gas. Disk 0 0 2 / 1 - 1 - 0 - 0 - 0 - 0 / 4
Orbital Mech. 0 0 3 / 3 - 20 - 0 - 0 - 2 - 0 / 22
GC FP 1 3 3 / 3 0.2469 16 - 1 - 0 - 0 - 0 / 19
GeV TeV ratio 4 8 30 / 11 0.0385 (0.7002) $\downarrow$ 1 - 0 - 0 - 0 - 0 / 1
Faber-Jackson 427 1947 1424 / 6 0.0101 (0.6445) $\uparrow$ 5 - 0 - 11 - 5 - 0 / 29
Diameter 3 3 6 / 2 0.0089 (0.0504) $\uparrow$ 18 - 0 - 2 - 0 - 0 / 31
D-Sigma 551 2901 1995 / 9 0.0 (0.0) 9 - 0 - 4 - 1 - 0 / 25
FP 967 1675 130214 / 22 0.0 (0.0) 10 - 0 - 5 - 2 - 0 / 28
IRAS 405 640 2947 / 2 0.0 (0.0) 4 - 1 - 16 - 0 - 0 / 35
Mass Model 2 4 7 / 7 0.0 (0.8942) $\downarrow$ 0 - 0 - 0 - 0 - 0 / 10
Tully-Fisher 9107 231663 56436 / 78 0.0 (0.0) 13 - 1 - 5 - 3 - 0 / 48
H I + optical distribution 0 0 1 / 1 - 0 - 0 - 0 - 0 - 0 / 1
LSB galaxies 0 0 2 / 1 - 7 - 0 - 0 - 1 - 0 / 10
Magnitude 0 0 111 / 3 - 13 - 0 - 5 - 12 - 0 / 40
Sosies 0 0 287 / 3 - 10 - 0 - 7 - 1 - 0 / 21
Tertiary 0 0 284 / 4 - 6 - 0 - 3 - 2 - 0 / 22
Tully est 0 0 1434 / 1 - 17 - 0 - 8 - 0 - 0 / 41
We are also interested in studying the tendency of each method to produce smaller or larger distances relative to the others. This will help to determine the methods that under- or overestimate distances with respect to other methods. For any given method pair and galaxy, we calculate the uncertainty-weighted mean distance modulus as
$\bar{\mu}^{i}_{g}=\frac{\sum_{k}\mu^{i}_{g,k}/(\sigma^{i}_{g,k})^{2}}{\sum_{k}1/(\sigma^{i}_{g,k})^{2}}$ (5)
and its uncertainty $\sigma^{i}_{g}=1/\sqrt{\sum_{k}{(\sigma^{i}_{g,k}})^{-2}}$ where $i$ represent the indicator, $g$ the galaxy and $k$ goes over all the measurements for this indicator and galaxy. We get these values for both methods. Then, we compute the uncertainty-weighted mean across all galaxies
$\bar{\mu}_{i}=\frac{\sum_{g}\bar{\mu}^{i}_{g}/(\sigma^{i}_{g})^{2}}{\sum_{g}1/(\sigma^{i}_{g})^{2}}$ (6)
and its uncertainty $\sigma_{\bar{\mu}_{i}}=1/\sqrt{\sum_{g}(\sigma^{i}_{g})^{-2}}$. We then construct a difference $X_{ij}\equiv\bar{\mu}_{i}-\bar{\mu}_{j}$, with uncertainty $\sigma_{X_{ij}}=\sqrt{\sigma_{\bar{\mu}_{i}}^{2}+\sigma^{2}_{\bar{\mu}_{j}}}$. This defines $X_{ij}$ for all methods with at least one galaxy in common, which we then use as constraints to infer the $\bar{\mu}_{i}$. Since not all method pairs have an $X_{ij}$ some methods will have more constraints than others, leading to better-determined $\bar{\mu}_{i}$. There is even the possibility of having methods with zero constraints or only one constraint. The former would imply that we cannot determine $\bar{\mu}_{i}$ for that method while the latter can cause degeneracies in the $\bar{\mu}_{i}$ of different indicators. It is important to realise that $\bar{\mu}_{i}$ represents an average relative distance inferred by method $i$, which must therefore be defined relative to an arbitrary zero-point. We take this to be the mean of all the $\bar{\mu}$. We constrain the $\bar{\mu}_{i}$ using the NUTS method of Hamiltonian Monte Carlo as implemented in the NumPyro sampler (Bingham et al., 2019; Phan et al., 2019), with wide uniform priors on the $\bar{\mu}_{i}$, ensuring convergence through the Gelman–Rubin statistic.
4 Results
4.1 NED-D
From the 76 indicators in NED-D, only 66 have at least one galaxy containing either two measurements with that indicator or one with that indicator and one with another. The others, which we cannot analyse for consistency and hence neglect for the remainder of the paper, are Quasar spectrum, SGRB, PAGB Stars, S Doradus Stars, GC K vs. (J-K), CO ring diameter, Radio Brightness, L(H{beta})-{sigma}, Dwarf Ellipticals and Jet Proper Motion.
To set the stage we show in Fig. 1 the distributions of distances probed by each method we consider. For each method the box shows the median distance (orange bar) and interquartile range, while the whisker covers the full range. The methods are ordered in median distance. We can see that 17 indicators (from SX Phe Stars to Mass Model) measure median distances lower than $0.1$ Mpc. Then, we have seven indicators (from Miras to GC SBF) having distances from 0.1 Mpc to 1 Mpc and nine from 1 Mpc to 10 Mpc (Wolf-Rayet to PNLF). These three ranges represent the first rung in the distance ladder. The second rung goes from a few Mpc to around several tens of Mpc (Wolf-Rayet to Black Hole). The third rung covers $\sim$100 Mpc to $\sim$400 Mpc. The SNIa method extends to thousands of Mpc, while GRB and HII LF reach even beyond 10 Gpc albeit with large relative uncertainties ($\gtrsim 30\%$). We also show in red the median fractional uncertainty of all measurements using a given indicator in NED-D, with scale on the right-hand y-axis. The methods highlighted in bold are those that we will find to be most consistent. Most of these have median relative errors lower than 20%, and two even below 10% (M Stars and BL Lac Luminosity). BL Lac Luminosity also probes to several Gpc. To be of maximum usefulness in the distance ladder, an indicator would ideally be self-consistent and consistent with other (good) methods, and probe large distances with small uncertainty.
Figure 1: Whisker plot showing the full range (whiskers) and interquartile range (boxes) of the distances probed by each indicator in NED-D. The median distances are shown by orange bars. We also show the median fractional distance uncertainty of each method in red, with scale on the right-hand y-axis.
Our consistency results are summarised in Tables 1 for the standard candle distance indicators and 2 for the standard ruler and secondary indicators. We report, for each method, the number of galaxies that have at least two distance measurements using the given indicator, the number of measurement pairs for these galaxies, the total number of measurements in the NED-D catalogue from this indicator and the number of references reporting these measurements, the KLD $p$-value (with a traffic-light-style colouring indicating the severity of the inconsistency) and external consistency statistics. For self-inconsistent cases we show the $p$-value after rescaling the uncertainties to give $\Delta_{ij}$ the variance of a standard half-normal distribution in parentheses. The final column contains the number of other distance indicators with which the method in question is consistent ($p>0.05$), in five different hyphen-separated cases. The first number describes the fiducial case with no shifting of means or rescaling of uncertainties. The second number gives the number of pairs that are made consistent when the means of the $\mu$ distributions are made to coincide, thus eliminating discrepant central values as a cause of inconsistency. The third number is the same but for rescaling the uncertainties so that the $\Delta_{ij}$ distribution has the width of a standard half-normal by construction, eliminating discrepant uncertainties as a cause of uncertainty. The fourth number corresponds to the case in which either shifting the means or rescaled the uncertainties produces consistency, and the fifth to the case where doing both is required to produce consistency. The final number is the total number of other methods with which that method has overlap, i.e. the largest number of pairs that could possibly be consistent.
Tables 1 and 2 show that most of the NED-D distance indicators are internally inconsistent, largely for an unidentified reason. This indicates that the catalogued distances are subject to problems like systematics and statistical outliers and are likely unsuitable for precision analyses such as constraining the Hubble constant. Indeed, the methods normally used for inferring $H_{0}$—Cepheids, TRGB and SNIa—are among the methods with the highest number of measurement pairs and all have a consistency $p$-value of 0. We also show the results for only the measurements that can be standardised, i.e. the ones with a quoted value for either $H_{0}$ or $\mu_{LMC}$. These results are shown in Appendix A (Figs. 4 and 5), where we see for example that the CMD method becomes much better behaved.
Figure 2: Consistency $p$-value results for the NED-D distance indicators. The diagonal squares describe internal consistency whereas the non-diagonal squares describe external consistency with other methods. We show the cases with $p$-value higher than 0.05 (i.e. not rejecting the null hypothesis) in green, the cases with no reported $p$-value in white and the cases with a $p$-value lower than 0.05 in a logarithmic colour scale. We show cases with underestimated uncertainties with a backslash hatching, overestimated uncertainties with forward-slash hatching, and inconsistent means with circle hatching. If either shifting the means or rescaling the uncertainties produces consistency we show both types of hatching, while if both are required we show star hatching. If the cause of the inconsistency is neither one of these nor their combination (i.e. it remains inconsistent after shifting the means and rescaling the uncertainties) we show no hatching. We call this an essential inconsistency.
Figure 3: Red: number of inconsistent method pairs (having $p<0.05$) as a function of the number of discarded or remaining methods, assuming that discarded methods are responsible for inconsistency in all comparisons involving them. Blue: number of fully consistent methods (having 100% external consistency) as a function of the number of discarded methods. We also show the total number of methods remaining on the top axis.
For internal inconsistency, we show with an arrow whether the inconsistency indicates under- or overestimated uncertainties. Some indicators present underestimated uncertainties (Horizontal Branch, AGN time lag, Type II Cepheids, G Lens, HII region diameter, GeV TeV ratio, and Mass Model) while Delta Scuti, Carbon Stars, Dwarf Galaxy Diameter, Ring Diameter, Faber-Jackson and Diameter are the opposite. For the remainder ones, the cause of the inconsistency was not the inconsistent uncertainties. Some distance indicators have external consistency results but no internal consistency $p$-value because they have at most one distance measurement to any given galaxy. One indicator (GRB) has an internal consistency result but no external consistency results because all of its measurements are to galaxies having no distances by any other method.
Figure 4: Mean values and $1\sigma$ C.L. uncertainties for the average distance modulus $\bar{\mu}$ returned by each indicator relative to the mean of all. Negative values correspond to distances underestimated on average relative to the mean of all indicators and vice versa.
In Fig. 2 we summarize the results for all distance indicators. We show the $p$-values both for internal consistency along the diagonal and external consistency on the off-diagonal. Cases where no consistency test is possible due to no overlap in measurements are shown as empty squares. Green indicates consistency ($p>0.05$), while $p$-values below 0.05 are shown by the colour-bar, with darker red indicating lower $p$. As we use 10,000 mocks the $p$-value is calculated in units of 0.0001; anything below this (i.e. no mock KLD as large as that of the data) is shown in dark red. The hatching indicates the cause of inconsistency. Over- and underestimated uncertainties are shown by forward and backward slanting diagonal lines respectively, while inconsistent means are shown by the circle pattern. Both types of hatching indicate that both problems are present to some degree, such that resolving either one produces consistency, and star hatching that both problems are present to a greater degree such that resolving both is required to remove the inconsistency. If neither shifting the means nor rescaling the uncertainties removes the inconsistency we label it “essential”, indicated by no hatching.
Figure 2 contains all the information for the consistency of the distance indicators. The total number of consistent pairs is 427 (41.8% of all pairs). 221 are inconsistent due to have uncertainties that are either too large or too small (mostly too small). Seven have inconsistent means, 109 are made consistent by matching either the means or the uncertainties, and five only by matching both. For problems that our tests cannot detect such as combinations of the previous ones along with e.g. outliers or non-Gaussian behaviour, we show no hatching. There are 163 such cases. Finally, 1994 pairs have no shared galaxies and hence no $p$-value. (Note that each pair is counted twice in Tables 1 and 2, so that the sums of the corresponding columns are twice the numbers quoted above.) Combining a priori consistent cases with those where we have identified the problem yields 769 cases, representing 82.5% of the total. Hence, even though most distance indicators are internally inconsistent, close to half of the cross-analyses are consistent, while only around a fifth are inconsistent due to an undiagnosed problem.
Figure 5: Top: Normalized probability distribution function (pdf) of the difference in units of $\Delta_{ij}$ for the SH0ES 2022 catalogue. We show the theoretical pdf of the standard half-normal for comparison. Bottom: Histogram for the mock KLD values generated under the null hypothesis of consistency compared to the KLD value of the real data.
Table 3: As Table 1 but for the SH0ES 2022 catalogue. In this case all distances use the Cepheid method so there is no notion of external consistency. N3982 and N0691 are shown in orange to indicate that they would fail a two-tailed consistency test (their $\Delta_{ij}$ values form a distribution more similar to a standard half-normal than would be expected from statistical noise under the assumption of consistency between all the measurements).
Name Measurement pairs KL $p$-value
Full dataset 345615 0.6613
N4258 97903 0.6896
LMC 115921 0.1286
N4038 406 0.9992
N1015 153 0.9808
N7678 120 0.953
N2525 2628 0.933
N1309 1378 0.8774
N7541 528 0.7556
U9391 528 0.7356
N4680 55 0.661
N0976 528 0.6108
N0691 378 0.553
N5584 13530 0.5328
N5728 190 0.5194
N1559 5995 0.464
N5861 820 0.4522
N3972 1326 0.4354
N5468 4278 0.4062
N4639 435 0.3974
N3254 1128 0.3972
N4536 780 0.3938
N3982 351 0.2766
N1365 1035 0.2418
N2608 231 0.1936
N3147 351 0.1834
M101 33411 0.1768
N7250 210 0.1756
N5917 91 0.149
N3447 5050 0.1276
N3021 120 0.1174
N1448 2628 0.117
Mrk 1337 105 0.1142
N3370 2628 0.0778
N2442 15576 0.042 (0.587) $\uparrow$
N3583 1431 0.0242 (0.2722) $\uparrow$
N5643 31375 0.0196 (0.851) $\uparrow$
N0105 28 0.0166 (0.4116) $\downarrow$
N7329 465 0.0154 (0.6016) $\downarrow$
M31 1485 0.015 (0.643) $\downarrow$
N4424 36 0.0004 (0.1314) $\downarrow$
In Fig. 3, we present the percentage of inconsistent method pairs and of fully externally consistent methods. This is shown as a function of the number of discarded or remaining methods (see Sec. 3.2), where it is assumed that each discarded method is responsible for inconsistency in any pair in which it is present. We continue discarding methods until we obtain the maximum set of fully externally consistent methods. We find this set to contain the following 18 methods: AGN time lag, PNLF, Sosies, M Stars, LSB galaxies, Proper Motion, Novae, SX Phe Stars, Maser, Dwarf Galaxy Diameter, GC FP, Subdwarf fitting, OB Stars, Orbital Mech. , B Stars, AGB, BL Lac Luminosity, GeV TeV ratio. It is interesting to compare this with the set of 12 self-consistent methods: BL Lac Luminosity, RSV Stars, OB Stars, Novae, Miras, RV Stars, Statistical, Brightest Stars, M Stars, Proper Motion, Maser, GC FP. The intersection of these sets gives us a set of 7 methods simultaneously achieving perfect internal and external consistency, which we highlight in bold throughout the paper: M Stars, Proper Motion, Novae, Maser, GC FP, OB Stars, BL Lac Luminosity. If we showed Fig. 2 with only these indicators, all the non-empty squares would be green.
The maximum set of fully externally consistent methods is not the same as the set of self-consistent methods. The set of methods being fully externally consistent but self-inconsistent contains 11 methods: AGN time lag, PNLF, Sosies, LSB Galaxies, SX Phe Stars, Dwarf Galaxy Diameter, Subdwarf fitting, Orbital Mech., B Stars, AGB, GeV TeV ratio. On the other hand, the set of methods being self-consistent but not fully externally consistent contains 5 methods: RSV Stars, Miras, RV Stars, Statistical, Brightest Stars. It is interesting to see that being fully externally consistent does not guarantee being internally consistent. This can be caused by a consistent subset of an indicator agreeing with other distance indicators or by both indicators being inconsistent in a way that cancels out when they are compared.
In Fig. 4, we present a whisker plot for the mean value and minimum $1\sigma$ C.L. uncertainties for the constraints on $\bar{\mu}_{i}$, a measure of the overall distance modulus returned by each method relative to the average of all of them. We present them in order for visual clarity. We discard two indicators (BL Lac Luminosity and GeV TeV ratio) as they both had only one constraint equation, which linked them (i.e. they each overlapped with only one other method in one galaxy, and that method for one of this pair was the other). This causes a perfect degeneracy between their $\bar{\mu}$ values, making the marginalised posteriors unbounded.
Fig. 4 shows the methods that tend to overestimate and underestimate distance modulus in average. Mass model is the indicator with the strongest underestimation having a difference $\Delta\bar{\mu}\equiv\bar{\mu}_{i}-\left<\bar{\mu}\right>\approx-0.6$ which corresponds to an underestimation of the luminosity distance by $\sim$20%. Proper Motion, Black Hole, HII region diameter, SNII radio, HII LF, and Ring Diameter have a difference of lower then -0.25 implying underestimation of the distance relative to the mean by more than 10%. Three indicators (SNIa SDSS, H I + Optical distribution and Wolf-Rayet) have a difference of around $1.5$ and higher, representing an overestimation of distance relative to the mean by a factor of around 2 or more. If we look for these indicators in figure 2, we can see that they are not consistent in any of the analyses. This is expected since they appear to overestimate distances by a huge margin. AGN time lag shows a difference higher than 0.5 implying an overestimation of distances higher than 25%. Out of the 7 indicators achieving full external consistency and internal consistency, only one seems to be significantly misestimating distances relative to the mean: Proper Motion. The remainder of the fully consistent indicators return values close to the mean (BL Lac Luminosity has no constraint due to the perfect degeneracy with GeV TeV ratio). Note that of course the mean may be biased by inconsistent methods, so returning distances compatible with the mean is not a reliable indicator of a method’s quality.
4.2 SH0ES
We now move to the analysis of the independent SH0ES 2022 catalogue. Fig. 5 presents two histograms for this catalogue. The upper panel shows the normalized pdf of the differences between the distance moduli ($\Delta_{ij}$) and the comparison to the standard half-normal. The lower panel shows the histogram of KLD values over the 10000 mock datasets compared to the value in the real data (vertical blue line). The full results are shown in Table 3, which (except for the first row describing the full catalogue) contains the galaxy name, number of measurement pairs and KLD $p$-value. The first two galaxies (N4258 and LMC) are in the first rung of the distance ladder and are used to calibrate the Cepheid PL relation by means of geometric distances to those galaxies.
Fig. 5 and the top row of Table 3 indicate that the SH0ES 2022 catalogue agrees well with the null hypothesis of statistical consistency. The two galaxies used for the first rung that calibrates the PL relation are also consistent with the null hypothesis. It is particularly important for these to be consistent, as inconsistency there would propagate into all the other distances and hence the constraint on e.g. $H_{0}$. From the 38 galaxies in the second rung (M31 does not host a SNIa but it is included in the SH0ES catalogue), 31 are consistent while the remaining 7 are inconsistent due to underestimated uncertainties. Note that most of the $p$-values only narrowly pass the threshold for rejection of the null hypothesis of consistency; had we take a tighter 0.01 threshold for the $p$-value, only one galaxy would have been considered inconsistent (N4424). This contrasts with NED-D where the great majority of inconsistent cases have $p<0.01$. Most of the inconsistent galaxies have $<$2000 pairs (apart from N2442 and N5643) and hence play a relatively small role in the consistency of the whole catalogue and the calibration of $M^{0}_{B}$ and hence $H_{0}$. Four of the seven inconsistent galaxies have underestimated uncertainties (N0105, N7329, M31 and N4424) whereas N2442, N3583 and N5643 have overestimated uncertainties. Interestingly, the inconsistent galaxies having more than 15000 measurement pairs have overestimated uncertainties, which has the effect of overestimating the uncertainty on $H_{0}$. We highlight two galaxies in lighter orange (N4038 and N1015) that would be inconsistent under a two-tailed test due to having $p>0.975$. This is due to the KLD being surprisingly low relative to the mock data distribution, indicating that the data conforms more closely to a standard half-normal than would be expected under the assumption of consistency given the uncertainties.
[…]
5 Discussion
5.1 Implications of the results
Our main result is that many of the distance measurements in NED-D are subject to significant problems, including “unknown unknowns” such as unaccounted-for systematics and outliers as well as the “known unknowns” of misestimated uncertainties and inconsistent central distance values between indicators. This makes it challenging currently to use for inference of cosmological parameters; the only previous attempt to do so yielded a 21% precision estimate of $H_{0}$ ($70\pm 15$ km/s/Mpc) from an ensemble of estimates and galaxies comparable to that studied here (Steer, 2020). This is likely sufficiently low precision not to be significantly affected by the inconsistencies we identify here. Despite our results, NED-D provides an excellent starting point to build a robust set of distance ladder measurements for cosmological inference. This would be similar to the approach of collaborations like SH0ES that construct and employ smaller, more carefully curated—and hence more self-consistent—extragalactic distance datasets.
Indeed, we find the Cepheids in the SH0ES 2022 catalogue not to exhibit significant problems. This supports the findings of Riess et al. (2022) that the dispersion in the Cepheid measurements agrees with the fiducial uncertainties without any unexplained variance. However, it is important to note that our test is not sensitive to constant shifts in the distance modulus across all Cepheids in a given galaxy. This would leave the distance modulus difference between the Cepheids, $\Delta_{ij}$, unchanged. This effect could in principle be present if the Wesenheit relative magnitude $m^{W}_{H}$ or the fiducial Wesenheit absolute magnitude $M^{W}_{H,1}$ have an unaccounted-for constant systematic error. Such a global shift to the distance modulus to each Cepheid could alter the Hubble diagram arbitrarily when combined with SNIa.
There is a small set of methods that exhibit both internal and external consistency (M-stars luminosity, Novae, Masers, Globular Cluster Fundamental Plane, O- and B-type Supergiants, BL Lac Luminosity and Proper Motion). However all of these have a small number of measurement pairs (no more than 127), and hence would not afford good constraints on the distance ladder—or cosmological parameters derived therefrom—in their own right. A worrying observation in this regard is that it is harder to acquire sufficient evidence to reject the null hypothesis when the dataset is smaller, suggesting that these methods may also become classified as inconsistent as more data on them is collected.
It is important to emphasize that statistically consistency is a necessary but not sufficient condition for a method to be useful in the distance ladder. Another necessary condition is to have a sufficient number of measurements to bring significant constraining power. For this reason, the seven methods in our fully consistent set are inferior to others having many more measurements like Cepheids, CMD, RR Lyrae, TRGB, Tully-Fisher and others. Future work must therefore either focus on statistically consistent subsets of these methods, or develop ways of measuring distances using the methods we identifying as having naturally higher consistency to more galaxies. An interesting indicator in this regard is Faber-Jackson, which has a very high number of measurement pairs (1947) and a p-value of 0.0101, near the threshold for consistency. We see from Table 1 that the inconsistency is due to overestimated uncertainties, probably the least pernicious cause. Another interesting case is CMD, which is self-consistent when only taking the measurements standardised through $H_{0}$ or $\mu_{LMC}$ (Appendix A) Moreover, this indicator has 3454 measurement pairs, almost double Faber-Jackson when restricting to standardisable measurements. It therefore shows promise as a useful method for constraining the distance ladder in the future.
Containing as it does measurements collected over 40 years, it is perhaps unsurprising that the overall consistency of NED-D is low. To investigate this further, we try splitting some of the better-known methods with many measurements (Cepheids, TRGB, SNIa, RR Lyrae, CMD and Tully-Fisher) into three bins of measurement year with roughly equal numbers of measurements in each bin and calculating the $p$-value for self-consistency within each bin. In this way, we can see if there are signs of an improving consistency over time. However, we find that for all indicators and bins the $p$-value remains, with the sole exception of RR Lyrae which has a $p$-value of 0.0008 for the intermediate bin. This indicates that the inconsistencies are not solely due to poorer measurements in the past or discrepancies between analysis choices in the past versus more recently.
Some of the inconsistency derives from non-independent measurements. For example, NED-D includes three TRGB distances to N6822 from Fusco et al. (2012), yet two of these differ only in the TRGB calibration, not the measurements themselves, and the third is an average of these two. This is simple a limitation with the way the literature was trawled to compile the catalogue: the abstract of the paper itself presents just one independent measurement. Another example comes from Rich et al. (2014), where the distances to 39 Cepheids in the galaxy NGC 6822 are provided. However, six of them were excluded from the analysis in the paper due to problems like phase shifting, crowding and deviations from the PL relation. Yet all 39 measurements are included in NED-D, potentially contributing significantly to inconsistency.
By excluding these problematic and non-independent measurements, we would get more information on the possibility of reducing the inconsistency across NED-D. One way to do this might be to automate a search over the literature using an Artificial Intelligence Large Language Model (LLM) as opposed to the human searchers employed in the construction of NED-D. This could allow a pre-cleaning of the data capable of removing unreliable cases. The LLM could also potentially be used to homogenise the calibrations of the methods by identifying the anchor used in a given paper, and its assumed distance, and then rescaling the measurements appropriately. If the inconsistencies could be reduced sufficiently, enormously more measurements—from a great deal more distance indicators—could be used to calibrate the distance ladder for cosmological inference and other applications.
Our results could potentially help explain the Hubble tension. For example, Fig. 4 indicates that Cepheid distances are lower than the mean of all methods. If distance were underestimated in SNIa host galaxies, then, by $H_{0}\approx cz/D$, the Hubble constant would be overestimated. TRGB likewise produce distances below the mean. Note however that the mean of all methods (most of which are inconsistent) is not a reliable indicator of the truth, and this simple accounting neglects the possibility of relative differences between the calibrator and cosmological first-rung samples.
To close this subsection we briefly describe the physics of the methods that we find to be most consistent, achieving both internal and external consistency. M-stars luminosity uses a relationship between the absolute magnitude of red dwarf stars and the temperature-independent spectral index. However, it requires low-resolution spectroscopy and thus it can only be applied in nearby galaxies, like the LMC (Schmidt-Kaler & Oestreicher, 1998). The Novae standard candle comes from a relation between the absolute magnitude at the maximum and the rate of decline in the flux known as the MMRD relation (Della Valle & Livio, 1994). For O- and B-type Supergiants stars, there exists a relation between their absolute magnitude and their spectral type and luminosity class. This method was notably applied to 30 Doradus nebula in the LMC (Walborn & Blades, 1997). BL Lac Object Luminosity is a method based on calibrating the absolute magnitude of BL Lacertae objects. It was found that the dispersion of this is small, although standardising it robustly does require a K-correction and evolution correction (Sbarufatti et al., 2005).
The Maser standard ruler is based on the dynamics of a water maser in an accretion disk surrounding a supermassive black hole. Using Very Long Baseline Interferometry (VLBI) it is possible to determine individual maser spots, the velocity of the masers and the shape of the disk. The physical size can then be deduced from Kepler’s laws and the angular size, affording a determination of the angular diameter distance (Humphreys et al., 2004; Reid et al., 2019). The Globular Cluster Fundamental Plane (GCFP) relates velocity dispersions, radii, and mean surface brightness of globular clusters to their absolute magnitude. This has been used to determine the distance to M31 (Strader et al., 2009). Finally, the Proper Motion method compares a galaxies’ apparent motion on the plane of the sky to its absolute motion calculated dynamically. This is also only applicable locally where proper motions are measurable (e.g. Lépine et al. 2011).
5.2 Comparison with the literature
Our NED-D analysis builds upon Singh (2021), where possible evidence for systematic underestimation of distance uncertainties was found. This could imply strong systematics on the late-time measurements of $H_{0}$, for example (Freedman et al., 2019a; Riess et al., 2022; Anand et al., 2022; Scolnic et al., 2023; Riess & Breuval, 2024), which might overestimate the tension with Planck. Despite using a different metric to quantify tension, and a considerably more sophisticated statistical framework, we agree with  Singh (2021) that most NED-D methods are inconsistent.
The NED-D catalogue has been used in various ways. For example, Cappellari et al. (2011) estimated uncertainties in the distances to galaxies in the ATLAS3D sample estimated from a Virgocentric infall peculiar velocity model by correlating them against the direct measurements in NED-D. This produced a sizeable 27% distance uncertainty for 692 galaxies with at least one NED-D distance measurement. Ohlson et al. (2023) used NED-D along with the Local Volume Galaxy catalogue (Karachentsev et al., 2013) to create a catalogue of galaxies within 50 Mpc with known distances, to study for example their nuclear X-ray activity.
NED-D has also been used to constrain fifth forces induced by modified gravity. Desmond et al. (2019) used the measurements of 51 galaxies with Cepheids and TRGBs from NED-D to set a bound on $\Delta G$, an effective change to Newton’s constant due to a long-range fifth force. This test assessed only relative consistency between Cepheid and TRGB distances as a function of $\Delta G$, although we see from Fig. 2 that these indicators are essentially inconsistent in NED-D. The consistency may be worse under modified gravity, but it is not good under General Relativity. Similarly, Desmond et al. (2021) used the TRGB distance to the LMC from NED-D as a constraint in inferring the gravitational constant in this dwarf galaxy from the dynamics of Cepheids in detached eclipsing binary systems. Each of these studies may be impacted by the inconsistencies we unearth here, and modified gravity constraints based on direct distance comparisons is a topic we will revisit in future work.
Our SH0ES analysis is complementary to others studying possible systematics and model misspecification in the Cepheid–SNIa determination of $H_{0}$. For example, the error bars on $H_{0}$ are impacted by outliers present in the data (Becker et al., 2015). This has been addressed in more recent work by removing the Cepheids that differ from $3.3\sigma$ in the global fit of the PL relation, which represent 1.2% of the catalogue (Riess et al., 2022). Another potential systematic is an environment-dependence on SNIa absolute magnitude, in the sense that SNIa in locally star-forming environments are dimmer than those in passive environments (Rigault et al., 2015, 2020). Another consideration is a possible difference between Cepheid observations and data reduction in the more nearby anchor galaxies (MW or LMC) relative to farther away ones like N4258. This can be studied by using only N4258 as the anchor (Kushnir & Sharon, 2024). This possibility seems less likely in light of our tests assessing consistency in individual galaxies, in particular the finding that both the LMC and N4258 have internally consistent Cepheid distances. The inferred value of $H_{0}$ can also be passband-dependent, with (Uddin et al., 2024) suggesting lower values in the B-band relative to the H-band used in SH0ES. Finally, the James Webb Space Telescope (JWST) has provided new insights on the possible systematics in the distance ladder. In particular, it has been claimed that JWST validates the HST distance measurements in N4258 and five SNIa hosts, rejecting at high significance the possibility of unrecognized crowding of Cepheid photometry growing with distance (Riess et al., 2024a). Indeed, some early samples of JWST Cepheid distances calibrated with only N4258 agree at with their HST counterparts (Jang et al., 2022; Riess et al., 2024b).
Nevertheless there is a plethora of possible problems with the Cepheid distance ladder (Owens et al., 2021, 2022; Madore & Freedman, 2024b; Owens et al., 2024; Madore & Freedman, 2024a), as well as results from the Chicago-Carnegie Hubble Program (CCHP) using new data from the James Webb Space Telescope (JWST) indicating that $H_{0}\approx 70$ km/s/Mpc with an insignificant tension with Planck (Freedman et al., 2024; Hoyt et al., 2025). This agrees with earlier CCHP results based on TRGB (Freedman et al., 2019a), and includes distances measured by the novel J-region Asymptotic Giant Branch (JAGB) method which boasts a number of advantages over Cepheids and TRGB (Madore & Freedman, 2020; Lee et al., 2024).
6 Conclusions
We study the consistency of 66 different methods for estimating redshift-independent distances from the NASA/IPAC Extragalactic Database of Distances (NED-D; Steer et al. 2016) and SH0ES 2022 (Riess et al., 2022) catalogues. This is achieved by selecting galaxies that have at least two distance measurements by a given method (“internal consistency”) or at least one measurement by each of two different methods (“external consistency”). We use the Mahalanobis distance to compute the distance differences in units of sigma, forming a distribution for each consistency test. If the distance estimates are consistent (scattered from a common value by their quoted uncertainties), this distribution should form a standard half-normal. To test if this is the case, we compute the Kullback–Leibler (KL) divergence of the empirical distribution and of 10,000 realizations of random samples under the null hypothesis of consistency and according to the measurements’ uncertainties. We compute the $p$-value as the fraction of realisations with a KL divergence higher than the value in the real data, adopting a threshold of 0.05 to discard the null hypothesis of consistency. This indicates significant evidence for under- or overestimated uncertainties, inconsistent means, statistical outliers, non-Gaussian uncertainties and/or systematic errors. When $p>0.05$, there is not enough evidence to discard the null hypothesis, which we denote by “consistent.” We perform a similar analysis for the SH0ES 2022 catalogue.
For NED-D we find that only 12 distance indicators achieve internal consistency. The maximal set of perfectly externally consistent methods has size 18 (i.e. all methods in this set are consistent with each other). The intersection of both sets gives the seven most reliable methods, which are both internally and externally consistent; these are M Stars luminosity, Proper Motion, Novae, Maser, Globular Cluster Fundamental Plane, O- and B-type Supergiants and BL Lac Object Luminosity. However, their number of measurements with these methods is low, generally $\mathcal{O}(10)$. This makes them unsuitable for a reconstruction of the distance ladder as they currently stand. A necessary condition besides statistical consistency for usefulness in the distance ladder is having a high number of measurements, and thus a better approach than focusing on these methods may be to build a consistent subset of distances using the indicators that provide many measurements. This paper gives a starting point for such constructions.
We also calculate the average measured distance modulus of each indicator across all galaxies relative to the mean of all indicators. We find that from the best set of indicators, Proper Motion is the only one with a considerable discrepancy in average returned distance (17.1% underestimation) relative to the mean. Outside the set of best indicators, some overestimate the luminosity distances by more than 95%; these are SNIa SDSS, H I + optical distribution and Wolf-Rayet. AGN time lag shows an overestimation higher than 25% and Mass model an underestimation higher than 20%. This indicates that some distance methods in NED-D are strongly affected by systematics and thus unsuitable for cosmological studies. We perform 932 analyses in total, including both internal and external consistency, of which 427 are consistent (41.8% of the total and 12 of them are from internally consistency analyses and the remainder externally consistency analyses). We repeat the analyses for the inconsistency cases by shifting the means and rescaling the uncertainties and find that 342 cases are made consistent by one or more of these modifications. In the remaining $\sim$1/5 of cases the inconsistency is not due only to incorrect central values or uncertainties, and we cannot identify the cause.
The SH0ES 2022 catalogue exhibits better behaviour. The KL divergence of the full catalogue yields a $p$-value of 0.66, in agreement with the null hypothesis of consistency. There is thus no evidence for hidden systematics in the Cepheid distances of this catalogue. We repeat this analysis for each of the 40 galaxies in the catalogue separately. The two galaxies in the first rung (N4258 and LMC) that calibrate the PL relation have consistent $p$-values as well (0.69 and 0.13). This is crucial because an inconsistency in the first rung would propagate to the second and third rungs, potentially biasing the entire distance ladder. From the second-rung galaxies, we find that 7 out of the 38 present an inconsistency ($p<0.05$) in a one-tailed test, and 8 in a two-tailed test. These galaxies however play a relatively minor role in the calibration of the SNIa absolute magnitude for which the catalogue was designed. We show this by repeating the MCMC excluding these inconsistent galaxies, finding very similar $H_{0}$.
In summary, the NED-D distance indicator measurements are subject to problems such as systematics, outliers, non-Gaussian behaviour and misestimated uncertainties. This is perhaps unsurprising given the heterogeneity of the measurements included, going back $\sim$45 years. Thus the full catalogue is not suitable for construction of a robust distance ladder. Even so, several papers have used the NED-D data for studies in cosmology (Freedman et al., 2019b; Chaparro-Molano et al., 2019; Steer, 2020; Zaninetti, 2023). By selecting a reliable subset of distances, it is however possible to determine galaxies’ distances accurately. Some of the inconsistency comes from non-independent and unreliable distance measurements that were impossible to exclude in the search procedure used to construct the catalogue. By removing these problematic data, the overall consistency might be considerably enhanced, enabling cosmological analysis; we will report elsewhere on our attempts in this direction. The more homogeneous and quality-controlled SH0ES 2022 catalogue, however, has no significant evident problems. The framework we develop may be simply applied to any catalogue of distances (or other quantities) to assess their consistency; we therefore hope that it will contribute to the future construction of a multiply cross-checked and fully robust distance ladder employing a wide range of methods.
Data availability
The code required to reproduce the results in this article are available at https://github.com/antonionajeraq/consistency_redshift_independent_galaxies.git. The NED-D data is available at https://ned.ipac.caltech.edu/Library/Distances/, and the SH0ES 2022 data at https://github.com/PantheonPlusSH0ES/DataRelease.
Acknowledgements
We thank Barry Madore, Claudia Maraston, Ian Steer, Richard Stiskalek and Tariq Yasin for useful inputs and discussion. JAN and HD are supported by a Royal Society University Research Fellowship (grant no. 211046).
B Stars 0 0 2 / 2 - 2 / 3
Black Hole 0 0 17 / 2 - 2 / 2
Blue Supergiant 0 0 2 / 1 - 3 / 5
BCG 0 0 239 / 1 - 2 / 3
Brightest Stars 0 0 129 / 41 - 0 / 6
Carbon Stars 0 0 82 / 13 - 12 / 16
Delta Scuti 0 0 12 / 2 - 4 / 7
FGLR 0 0 7 / 7 - 0 / 3
GCLF 0 0 781 / 96 - 5 / 16
GC SBF 0 0 3 / 1 - 6 / 9
HII LF 0 0 17 / 2 - 0 / 4
Horizontal Branch 0 0 110 / 53 - 5 / 11
M Stars 0 0 6 / 5 - 2 / 6
Miras 0 0 36 / 27 - 8 / 15
Novae 0 0 18 / 11 - 11 / 17
OB Stars 0 0 4 / 3 - 6 / 7
PNLF 0 0 183 / 57 - 2 / 6
Red Clump 0 0 236 / 63 - 8 / 14
RSV Stars 0 0 9 / 6 - 3 / 8
SX Phe Stars 0 0 3 / 2 - 0 / 1
Statistical 0 0 292 / 27 - 14 / 37
Subdwarf fitting 0 0 1 / 1 - 0 / 1
SZ effect 0 0 283 / 19 - 4 / 5
SBF 0 0 1781 / 67 - 2 / 14
SNII radio 0 0 24 / 2 - 2 / 6
White Dwarfs 0 0 1 / 1 - 0 / 1
Grav. Wave 0 0 8 / 8 - 1 / 1
Appendix A Results with Standardisable Measurements Only
Here we show the analogues of Tables 1 and 2 but excluding measurements that do not list a $H_{0}$ or $\mu_{\text{LMC}}$ value and hence cannot be standardised. On the whole we find that the consistency little improves, indicating that the non-standardisability of many of the measurements was not a principal cause of inconsistency in the main text. For this reason we do not show the results after rescaling the uncertainties and/or shifting the means. We do however find that behaviour of the CMD indicator improves dramatically. This method also has a significant number of measurements, suggesting that it may be useful for constraining the distance ladder.
Table 5: As Table 2 but including only measurements that can be standardised through either $H_{0}$ or $\mu_{\text{LMC}}$.
Distance indicator # of galaxies # pairs # total measurements/references KL $p$-value External consistency
SNII optical 128 390 1622 / 89 0.0 1 / 7
Dwarf Galaxy Diameter 0 0 16 / 1 - 6 / 9
Eclipsing Binary 0 0 273 / 48 - 9 / 15
GC radius 0 0 101 / 8 - 1 / 11
G Lens 0 0 86 / 14 - 0 / 1
HII region diameter 0 0 53 / 7 - 1 / 12
Maser 0 0 24 / 20 - 4 / 8
Orbital Mech. 0 0 3 / 3 - 0 / 2
Proper Motion 0 0 5 / 3 - 2 / 3
Ring Diameter 0 0 212 / 1 - 2 / 4
GeV TeV ratio 3 5 30 / 11 0.0124 -
Faber-Jackson 427 1907 1424 / 6 0.0016 4 / 18
D-Sigma 548 2787 1995 / 9 0.0 9 / 16
FP 807 1134 130214 / 22 0.0 7 / 22
IRAS 404 639 2947 / 2 0.0 3 / 28
Tully-Fisher 4972 99919 56436 / 78 0.0 11 / 44
Diameter 0 0 6 / 2 - 9 / 15
GC FP 0 0 3 / 3 - 7 / 9
H I + optical distribution 0 0 1 / 1 - 0 / 1
LSB galaxies 0 0 2 / 1 - 1 / 4
Magnitude 0 0 111 / 3 - 1 / 9
Sosies 0 0 287 / 3 - 1 / 4
Tertiary 0 0 284 / 4 - 5 / 12
Tully est 0 0 1434 / 1 - 17 / 27